# Enhancing Visual Place Recognition via Fast and Slow Adaptive Biasing in Event Cameras

### v1

@@ Abstract
Event cameras are increasingly popular in robotics due to their beneficial features, such as low latency, energy efficiency, and high dynamic range. Nevertheless, their downstream task performance is greatly influenced by the optimization of bias parameters. These parameters, for instance, regulate the necessary change in light intensity to trigger an event, which in turn depends on factors such as the environment lighting and camera motion. This paper introduces feedback control algorithms that automatically tune the bias parameters through two interacting methods: 1) An immediate, on-the-fly fast adaptation of the refractory period, which sets the minimum interval between consecutive events, and 2) if the event rate exceeds the specified bounds even after changing the refractory period repeatedly, the controller adapts the pixel bandwidth and event thresholds, which stabilizes after a short period of noise events across all pixels (slow adaptation). Our evaluation focuses on the visual place recognition task, where incoming query images are compared to a given reference database. We conducted comprehensive evaluations of our algorithms’ adaptive feedback control in real-time. To do so, we collected the QCR-Fast-and-Slow dataset that contains DAVIS346 event camera streams from 366 repeated traversals of a Scout Mini robot navigating through a 100 meter long indoor lab setting (totaling over 35km distance traveled) in varying brightness conditions with ground truth location information. Our proposed feedback controllers result in superior performance when compared to the standard bias settings and prior feedback control methods. Our findings also detail the impact of bias adjustments on task performance and feature ablation studies on the fast and slow adaptation mechanisms.
@@ I Introduction
Event cameras distinguish themselves from traditional cameras by their unique operational principle. Unlike traditional cameras, which capture images at set intervals, event cameras generate data when individual pixels detect changes in brightness—leading to either an OFF or an ON event when the change falls below or exceeds a certain threshold. The resulting output is an event stream that includes the $(x,\ y)$ pixel coordinates, a precise timestamp $t$ and the polarity $p$ indicating the event type (OFF/ON). This sparse and asynchronous data capture method grants event cameras high temporal resolution, low latencies, and reduced power consumption; qualities that could be advantageous for autonomous systems, particularly in tasks like motion segmentation, pose estimation, visual odometry, simultaneous localization and mapping (SLAM) and neuromorphic control [1].
@@
We demonstrate our proposed feedback controllers on the Visual Place Recognition (VPR) task [2, 3, 4, 5]. VPR benefits the long-term autonomy of mobile robots as a core component in localization and mapping algorithms, including loop-closure in SLAM and for global re-localization in kidnapped robot scenarios. Event-based VPR has seen an increased interest over the past few years [6, 7, 8, 9, 10, 11, 12, 13], but an in-depth investigation of the impact of the bias parameters on VPR performance is lacking until now.
Figure 1: Closed-loop biasing of event cameras. We perform feedback control on event camera biases to obtain improved performance for the visual place recognition (VPR) downstream task. We achieve this by varying refractory period, pixel bandwidth, and event threshold parameters in real-time. As illustrated, we modulate these device parameters to monitor and maintain event-rate from an onboard event-camera mounted on a mobile robot.
@@
To support ongoing research in this area, we will release both the algorithm and the QCR-Fast-and-Slow Dataset [^1] We will provide the download link in the final version..
@@ II-A Event stream representations and processing
When processing event streams in real-time, the high temporal resolution requires algorithms that can process events rapidly. It is desirable to keep the event rate within certain bounds, as higher event-rates can lead to processing delays. [19] maintains desirable event-rates through various event-dropping strategies at the hardware level. [20] tackles process overflow due to high rates of event generation by modulating event package sizes depending on expectations of process times and hardware limitations, leading to a graceful degradation of process output through controlled discarding of events. [21] utilizes a gain parameter to adapt its processing capabilities to ensure the absolute delay of the system is controlled.
Tackling process overload through controlled discarding of events leads to task-relevant data loss. [22] adapts ON/OFF event thresholds in the device hardware to monitor and module event rates to prevent data losses. The popular jAER [23] exposes bias parameters along with underlying coarse and fine bias settings [24]. This opens up pathways to software-triggered adaptive biasing approaches, which couples various information sources to infer environmental conditions and tune the device on the fly.
The most relevant prior work is [14]. They perform feedback control to regulate event-rates and noise by changing event threshold, pixel bandwidth, and refractory period. Section IV-D contains a detailed explanation of the various feedback controllers introduced in [14]. [25] evaluates device output with changes in device biases on a Prophesee event camera. [26] characterizes DVS performances to a set of quantifiable parameters and evaluates DVS responses based on parameters such as threshold, bandwidth, refractory period and scene illumination.
@@ II-B Event-based Visual Place Recognition
[8] combines temporal windows of varying lengths through parallelized ensembles to obtain a high-performing VPR system at the cost of high computational requirements. [9] uses the popular Vector of Locally Aggregated Descriptors (VLAD) pooling mechanism to obtain image descriptors after reconstructing event edges. This idea has been taken further in [7] which introduced the first end-to-end deep-learning event-based VPR framework. As opposed to using conventional deep networks, [10] has introduced a spiking neural network to process the event stream. [6] introduced a computationally efficient technique that relies on just a small subset of informative pixels. Another interesting approach is the combination of conventional frames with event streams for improved performance [13].
Moving towards more bio-inspired approaches, [11] developed a brain-inspired multimodal hybrid neural network that seeks to imitate the neurological processes of the brain for more efficient robot place recognition. Their model demonstrates the synergy between different sensory modalities and computational paradigms, offering insights into the development of advanced robotic systems. This approach is similar to [10] who also used a spiking neural network for processing the event stream. Similarly, [12] harnessed the power of neuromorphic computing, presenting a neuromorphic sequence learning model that utilizes event cameras to navigate through vegetative environments.
@@ III-A Preliminaries: Event Camera Principles
Post-processing techniques exist to mitigate background noise [35, 34], but on-chip bias generators also play a crucial role in adapting the camera to varying lighting conditions [36, 37, 24]. As detailed in the next section, our methodology automates the adjustment of these bias parameters, akin to tuning converter gain, exposure time, and ISO in traditional cameras.
@@ III-B Fast changes in refractory period
Then, the event rate $R(t,\tau)$ at time $t$ over the time window of duration $\tau$ is defined as:
@@
Further, we ensure the bias settings remain within the desired bounds [$B_{\text{r}}^{-},B_{\text{r}}^{+}]$:
$B_{\text{r}}=\begin{cases}B_{\text{r}}^{+}&\textit{if }B_{\text{r}}>B_{\text{r}}^{+}\\ B_{\text{r}}^{-}&\textit{if }B_{\text{r}}<B_{\text{r}}^{-}.\end{cases}$ (4)
@@ III-C Slow changes in pixel bandwidth and event threshold
We vary the pixel bandwidth and event thresholds to maintain event rates within desirable range $R\in[R^{-},R^{+}]$. However, the instances of these changes have to be monitored because changing pixel bandwidth or event thresholds lead to momentarily disruptive event bursts in the output. Hence slow and simultaneous fixed-step adaptations $\Delta B_{\text{p}}$ and $\Delta B_{\text{e}}$ are performed in pixel bandwidth $B_{\text{p}}$ and event thresholds $B_{\text{e}}$ respectively if the event rate remains outside the desired bounds for the previous $N$ consecutive time samples. Specifically, at time $t$:
@@ IV Experimental setup
Figure 6: (Top) Robot System Setup. Left: Scout Mini platform equipped with various sensors, including a LiDAR to obtain pseudo ground truth and the DAVIS346 event camera mounted forward-facing. Right: Close-up of the DAVIS346 event camera, where the aperture setting can be manually adjusted. Altering the aperture simulates different brightness conditions: closing it simulates low-light environments, while opening it simulates high-brightness scenarios. (Bottom) Map from the robot. The blue colored segment marks the starting position and red arrows indicate the traversal path.
@@ IV-A Robot platform
We mount an iniVation DAVIS346 event camera [40] in the front of an AgileX Scout Mini mobile robot such that the camera faces in the direction of motion (Figure 6, left). We ensure that the camera focus remains constant in all our experiments, and utilize manual changes in camera aperture to simulate various environment brightness conditions (Figure 6, right).
@@ IV-B Closed-loop bias controller
Events are streamed over UDP for real-time computation of event rates. Our feedback bias controller utilizes event rates as feedback to control pixel bandwidth, event thresholds and refractory period in an online manner. Following up on Equation 1, the event rates are calculated and updated over time intervals of $\tau=300ms$.
@@
IV-B1 Bias initializations
@@ IV-C Visual Place Recognition Pipeline
We use a previously proposed [43], simple visual place recognition pipeline to demonstrate our feedback controller. To ensure robustness against appearance changes, the event frames are patch-normalized [44] with a patch-size of $8px$. Sum-of-absolute-differences (SAD) is performed between these event frames to obtain distance matrices – the higher the SAD, the less similar the two event frames are [44]. The event frame in the reference dataset that has the lowest SAD to the current query event frame is considered as the top match.
@@ IV-D4 Feedback control of event threshold to bound noise
The event threshold determines how much of a change is required in log input light intensity for an event spike to occur. The fourth baseline [14] exploits event thresholds to maintain event rates within desirable bounds. Unlike refractory period control, changes in the event threshold lead to event bursts, making the device’s output momentarily noisy. In contrast, similar to Section IV-D3, our slow adaptation approach performs sparse and simultaneous changes in pixel bandwidth and event threshold parameters while monitoring the event rate.
@@ V-A Comparison to state-of-the-art
The same trend can be observed in Table II (reference dataset: low brightness), where the performance improvement over the second-best-performing pixel bandwidth controller [14] is an absolute 4% in the high brightness query condition (i.e. the query condition with the highest appearance change). Figure 8 demonstrates that the performance advantage in the worst-case run becomes even more pronounced, comparing a R@1 of 85% of our technique with 59% (default parameters), 73% (pixel bandwidth controller), 59% (refractory period controller) and 56% (pixel threshold controller).
@@ V-B2 Components
Figure 9 (right) shows the R@1 performance from three iterations of each variation. When the query set is recorded in low brightness conditions, all variants significantly improve upon the constant baseline technique. In the medium and high brightness query conditions, the performance differences are relatively small; with our fast+slow controller achieving the highest performance in both of these conditions. Interestingly, only changing the refractory period (‘fast’) leads to a small performance decrease when compared to the constant baseline.
@@ VI Discussion and Conclusions
This paper has introduced a novel feedback control approach for event cameras, integrating fast-and-slow mechanisms to dynamically adjust bias parameters, thereby optimizing event stream processing for VPR tasks. Our methodology not only enhances the adaptability of event cameras to varying environmental conditions but also demonstrates substantial improvements in place recognition performance.
@@
Looking ahead, several avenues for further research emerge from this study. A promising direction is the integration of our feedback control system with Spiking Neural Networks (SNNs) [45, 46, 47, 10, 11, 12]. SNNs, with their bio-inspired processing capabilities, could significantly benefit from the refined event streams produced by our method, potentially leading to more robust and efficient visual recognition systems.

### v2

# PREPRINT VERSION | PAPER ACCEPTED TO THE IEEE/RSJ INTERNATIONAL CONFERENCE ON INTELLIGENT ROBOTS AND SYSTEMS (IROS 2024)

@@ Abstract
Event cameras are increasingly popular in robotics due to beneficial features such as low latency, energy efficiency, and high dynamic range. Nevertheless, their downstream task performance is greatly influenced by the optimization of bias parameters. These parameters, for instance, regulate the necessary change in light intensity to trigger an event, which in turn depends on factors such as the environment lighting and camera motion. This paper introduces feedback control algorithms that automatically tune the bias parameters through two interacting methods: 1) An immediate, on-the-fly fast adaptation of the refractory period, which sets the minimum interval between consecutive events, and 2) if the event rate exceeds the specified bounds even after changing the refractory period repeatedly, the controller adapts the pixel bandwidth and event thresholds, which stabilizes after a short period of noise events across all pixels (slow adaptation). Our evaluation focuses on the visual place recognition task, where incoming query images are compared to a given reference database. We conducted comprehensive evaluations of our algorithms’ adaptive feedback control in real-time. To do so, we collected the QCR-Fast-and-Slow dataset that contains DAVIS346 event camera streams from 366 repeated traversals of a Scout Mini robot navigating through a 100 meter long indoor lab setting (totaling over 35km distance traveled) in varying brightness conditions with ground truth location information. Our proposed feedback controllers result in superior performance when compared to the standard bias settings and prior feedback control methods. Our findings also detail the impact of bias adjustments on task performance and feature ablation studies on the fast and slow adaptation mechanisms.
@@ I Introduction
Event cameras distinguish themselves from traditional cameras by their unique operational principle. Unlike traditional cameras, which capture images at set intervals, event cameras generate data when individual pixels detect changes in brightness, leading to OFF or ON events when the change falls below or exceeds a certain threshold. The resulting output is an event stream that includes the $(x,\,y)$ pixel coordinates, a high-precision timestamp $t$, and the polarity $p$ indicating the event type (OFF/ON). This sparse and asynchronous data capture method grants event cameras high temporal resolution, low latencies, and reduced power consumption; qualities that could be advantageous for autonomous systems, particularly in tasks like motion segmentation, pose estimation, visual odometry, simultaneous localization and mapping (SLAM) and neuromorphic control [1].
@@
We demonstrate our proposed feedback controller on the Visual Place Recognition (VPR) task [2, 3]. VPR benefits the long-term autonomy of mobile robots as a core component in localization and mapping algorithms, aiding loop-closure in SLAM and global re-localization in kidnapped robot scenarios. Event-based VPR has seen an increased interest over the past few years [4, 5, 6, 7, 8, 9, 10, 11], but an in-depth investigation of the impact of the bias parameters on VPR performance is lacking until now.
Figure 1: Closed-loop biasing of event cameras. We perform feedback control on event camera biases to obtain improved performance for the visual place recognition downstream task. We achieve this by varying refractory period, pixel bandwidth, and event threshold parameters in real-time. As illustrated, we modulate these device parameters to monitor and maintain event-rate from an onboard event-camera mounted on a mobile robot.
@@
To support ongoing research in this area, we release both the algorithm and the QCR-Fast-and-Slow Dataset: https://github.com/gokulbnr/fast-slow-biased-event-vpr
@@ II-A Event stream representations and processing
When processing event streams in real-time, the high temporal resolution requires algorithms that can process events rapidly. It is desirable to keep the event rate within certain bounds, as higher event-rates can lead to processing delays. [17] maintains desirable event-rates through various event-dropping strategies at the hardware level. [18] tackles process overflow due to high event rates by modulating event package sizes and controlled discarding of events depending on expectations of process times and hardware limitations. [19] uses a gain parameter to control the absolute delay of the system.
Tackling process overload through controlled discarding of events leads to task-relevant data loss. [20] adapts ON/OFF event thresholds in the device hardware to monitor and modulate event rates. The popular jAER software [21] exposes bias parameters along with underlying coarse and fine bias settings [22]. This opens up pathways to software-triggered adaptive biasing approaches, which couples various information sources to infer environmental conditions and tune the device on the fly.
The most relevant prior work is [12]. They perform feedback control to regulate event-rates and noise by changing event threshold, pixel bandwidth, and refractory period. Section IV-D contains a detailed explanation of the various feedback controllers introduced in [12]. [23] evaluates how changes in device biases affect the output of a Prophesee event camera, using sharpness to assess event stream quality. [24] characterizes DVS performances to a set of quantifiable parameters and evaluates DVS responses based on parameters such as threshold, bandwidth, refractory period and scene illumination.
@@ II-B Event-based Visual Place Recognition
[6] combines temporal windows of varying lengths through parallelized ensembles to obtain a high-performing VPR system at the cost of high computational requirements. [7] uses the popular Vector of Locally Aggregated Descriptors (VLAD) pooling mechanism to obtain image descriptors after reconstructing event edges. This idea has been taken further in [5] which introduced the first end-to-end deep-learning event-based VPR framework. [4] introduced a computationally efficient technique that relies on just a small subset of informative pixels. Another interesting approach is the combination of conventional frames with event streams for improved performance [11].
Moving towards more bio-inspired approaches, [9] developed a brain-inspired multimodal hybrid neural network that seeks to imitate the neurological processes of the brain for more efficient robot place recognition. Their model demonstrates the synergy between different sensory modalities and computational paradigms, offering insights into the development of advanced robotic systems. This approach is similar to [8] who also used a spiking neural network for processing the event stream. Similarly, [10] harnessed the power of neuromorphic computing, presenting a neuromorphic sequence learning model that uses event cameras to navigate through vegetative environments.
@@ III-A Preliminaries: Event Camera Principles
Post-processing techniques exist to mitigate background noise [33, 32], but on-chip bias current generators also play a crucial role in adapting the camera to varying lighting conditions [34, 35, 22]. As detailed in the next section, our methodology automates the adjustment of these bias parameters, akin to tuning converter gain, exposure time, and ISO in traditional cameras.
@@ III-B Fast changes in refractory period
Then, the event rate $R(t,\tau)$ (measured in events per second) at time $t$ over the duration $\tau$ is defined as:
@@
Further, we ensure the bias settings remain within the desired bounds [$B_{\text{r}}^{-},B_{\text{r}}^{+}]$, resulting in the final bias setting $\hat{B}_{\text{r}}$:
$\hat{B}_{\text{r}}=\begin{cases}B_{\text{r}}^{+}&\text{if }B_{\text{r}}>B_{\text{r}}^{+}\\ B_{\text{r}}^{-}&\text{if }B_{\text{r}}<B_{\text{r}}^{-}\\ B_{\text{r}}&\text{otherwise}.\end{cases}$ (4)
@@ III-C Slow changes in pixel bandwidth and event threshold
We vary the pixel bandwidth and event thresholds to maintain event rates within desirable range $R\in[R^{-},R^{+}]$. However, the instances of these changes have to be monitored because changing pixel bandwidth or event thresholds lead to momentarily disruptive event bursts in the output. Hence slow and simultaneous fixed-step adaptations $\Delta B_{\text{p}}$ and $\Delta B_{\text{e}}$ are performed in pixel bandwidth $B_{\text{p}}$ and event thresholds $B_{\text{e}}$ respectively if the event rate remains outside the desired bounds for the previous $N$ consecutive time samples despite fast changes in the refractory period (Section III-B). Specifically, at time $t$:
@@ IV Experimental setup
Figure 6: (Top) Robot system setup. Left: Scout Mini platform equipped with various sensors, including a LiDAR to obtain pseudo ground truth and the DAVIS346 event camera mounted forward-facing. Right: Close-up of the DAVIS346 event camera, where the aperture setting can be manually adjusted. Altering the aperture simulates different brightness conditions: closing it simulates low-light environments, while opening it simulates high-brightness scenarios. (Bottom) Map from the robot and frames from the event-data. Left: The blue colored segment marks the starting position and red arrows indicate the traversal path. Right: Frames of accumulated event-data taken from one of the traversals.
@@ IV-A Robot platform
We mount an iniVation DAVIS346 event camera [38] in the front of an AgileX Scout Mini mobile robot such that the camera faces in the direction of motion (Figure 6, left). We ensure that the camera focus remains constant in all our experiments, and use manual changes in camera aperture to simulate various environment brightness conditions (Figure 6, right). While no deliberate changes in the light source positions were introduced, they might have occurred naturally as the dataset was collected over a prolonged time. Although Active Pixel Sensor (APS) grayscale images and Inertial Measurement Unit (IMU) data were also recorded along with the event data, they were not used in the experiments.
@@ IV-B Closed-loop bias controller
Events are streamed over UDP for real-time computation of event rates. Our feedback bias controller uses event rates as feedback to control pixel bandwidth, event thresholds and refractory period in an online manner. Following up on Equation 1, the event rates are calculated and updated over time intervals of $\tau=300ms$.
@@
IV-B1 Bias current initializations (in Amperes)
@@ IV-C Visual Place Recognition Pipeline
We use a previously proposed, simple visual place recognition pipeline to demonstrate our feedback controller [41]. To ensure robustness against appearance changes, the event frames are patch-normalized [42] with a patch-size of $8px$. Sum-of-absolute-differences (SAD) is performed between event frames to obtain distance matrices – the higher the SAD, the less similar the two event frames [42]. The event frame in the reference dataset that has the lowest SAD to the current query event frame is considered as the top match.
@@ IV-D4 Feedback control of event threshold to bound noise
The event threshold determines how much of a change is required in log input light intensity for an event spike to occur. The fourth baseline [12] exploits event thresholds to maintain event rates within desirable bounds. Unlike refractory period control, changes in the event threshold and pixel bandwidth lead to event bursts, making the device’s output momentarily noisy. In contrast, similar to Section IV-D3, our slow adaptation approach performs sparse and simultaneous changes in pixel bandwidth and event threshold parameters while monitoring the event rate.
@@ V-A Comparison to state-of-the-art
The same trend can be observed in Table II (reference dataset: low brightness), where the performance improvement over the second-best-performing pixel bandwidth controller [12] is an absolute 4% in the high brightness query condition, the query condition with the highest appearance change. Figure 8 demonstrates that the performance advantage in the worst-case run becomes even more pronounced, comparing a R@1 of 85% of our technique with 59% (default parameters), 73% (pixel bandwidth controller), 59% (refractory period controller) and 56% (pixel threshold controller).
@@ V-B2 Components
Figure 9 (right) shows the R@1 performance from three iterations of each variation. When the query set is recorded in low brightness conditions, all variants significantly improve upon the constant baseline technique. In the medium and high brightness query conditions, the performance differences are relatively small; with our fast+slow controller achieving the highest performance in these conditions. Interestingly, only changing the refractory period (‘fast’) leads to a small performance decrease compared to the constant baseline.
@@ VI Discussion and Conclusions
This paper has introduced a novel feedback control approach for event cameras, integrating fast-and-slow mechanisms to dynamically adjust bias parameters, thereby optimizing event stream processing for VPR tasks. Our methodology not only enhances the adaptability of event cameras to varying environmental conditions but also demonstrates that a feedback controller built specifically for the downstream task of VPR outperforms a general feedback controller in the same downstream task.
@@
Future work will extend this approach to longer trajectories and larger environmental variations, such as day-night transitions and diverse weather conditions. We also plan to integrate our feedback control system with Spiking Neural Networks (SNNs) [43, 44, 8, 9, 10]. SNNs, with their bio-inspired processing capabilities, could significantly benefit from the refined event streams produced by our method, potentially leading to more robust and efficient visual recognition systems.